OpenStudy (hamzajani):

If f(g(x))=4x^2-8x and f(x) =x^2-4,then g(x)

TheSmartOne (thesmartone):

We plug in g(x) in to f(x) f(x) = x^2 - 4 f(g(x)) = 4x^2 - 8x You plug in g(x) in to f(x) to get f(g(x)) so basically f(x) = x^2 - 4 f(g(x)) = g(x)^2 - 4 However, we also know that f(g(x)) = 4x^2 - 8x so set each equal to each other and solve g(x)^2 - 4 = 4x^2 - 8x First add 4 on both sides and then take the square root on both sides. And ta-da!

OpenStudy (unklerhaukus):

\[f(g(x))=4x^2-8x\\\qquad\quad=(2x)^2-4(2x)\]